Consider the R^2 - R function defined by f(x,y) = 3x + 2y.

Prove from first principles that lim (x,y) -> (1,-1) f(x,y) = 1


1
Expert's answer
2022-05-24T17:06:18-0400

ANSWER

Let "\epsilon >0" , "\delta=\frac{\epsilon}{5}" and "\sqrt{(x-1)^{2}+(y+1 )^{2}}<\delta" . Since

"\left | x-1 \right |\leq \sqrt{(x-1)^{2}+(y+1 )^{2}}, \left | y+1 \right |\leq \sqrt{(x-1)^{2}+(y+1 )^{2}}" , then "|f(x,y)-1|=|3x+2y-1|=|(3x-3 )+(2y+2)|\leq 3|x-1|+2|y-(-1)|<\\<5\delta=5\cdot \frac {\epsilon}{5}=\epsilon."

So, "\lim _{(x,y)\rightarrow(1,-1)}(3x+2y)=1" .